What are the inverse of the functions f(x)=10^x and log10^x?
Answer by stanbon(26273)

What are the inverse of the functions
f(x)=10^x
---
Interchange x and y:
x = 10^y
Solve for "y":
y = log(x) is the inverse function.
==========================================
y = log10^x?
y = x
---
Interchange x and y:
x = y
Solve for "y":
---
y = x is the inverse function.
====================================
Cheers,
Stan H.